OpenStudy (anonymous):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 4 inches, and the height of the cone is 6 inches. If the diameter of the bubble gum is 0.8 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? Use 3.14 for pi. Select one: a. 81.67 cubic inches b. 91.23 cubic inches c. 92.81 cubic inches d. 100.21 cubic inches

OpenStudy (anonymous):

@Chris911

OpenStudy (anonymous):

@eliassaab help

OpenStudy (anonymous):

You are not going to fill the bubble gum with ice cream The volume of the cone \[ V= \frac 1 3 \pi r^2 h= \frac 13 (3.14) (4^2) 6=100.48 \]

OpenStudy (anonymous):

I will say d.